OpenStudy (anonymous):

3x-9=5x-7 WILL GIVE MEDAL FOR CORRECT ANSWER!

OpenStudy (anonymous):

Well you have the equation \[3x-9=5x-7\] which you then bring the x's ot one side and the numbers to the other \[3x-9+(9)=5x-7 +(9)\] \[3x-(5x)=5x-2-(5x)\] \[\frac{ -2x=-2 }{ -2 }\] \[x=1\]

OpenStudy (anonymous):

sorry its 9+(-7)=2 making x=-1